### v1

# A homogeneous geometry of low-rank tensors

Abstract
We consider sets of fixed CP, multilinear, and $T$ rank tensors, and derive conditions for when (the smooth parts of) these sets are smooth homogeneous manifolds. For CP and $T$ ranks, the conditions are essentially that the rank is sufficiently low. These homogeneous structures are then used to derive Riemannian metrics whose geodesics are both complete and efficient to compute.
1 Introduction
Identifying fixed-rank matrices as quotient manifolds has had very profitable applications in manifold optimization and statistics [Bonnabel10, Journee10, Bonnabel12, Mishra14, Massart20, Zheng25]. For example, Vandereycken, Absil, and Vandewalle’s [Vandereycken12] identification of fixed-rank symmetric matrices as a homogeneous manifold is particularly useful because it induces a Riemannian geometry with complete geodesics. A natural question is whether such a construction generalizes to tensors.
Consider the group action that multiplies tensors in each mode by an invertible matrix, which can also be seen as a change of basis in each mode. While there are several ways to define the rank of a tensor, for almost all of them, the rank is invariant under this action. There are in general two obstacles to using it to identify sets of fixed-rank tensors as homogeneous manifolds:
  1. 1.
    Without modification, the action is not transitive since there is no way to go from tensors on the form $a\otimes a\otimes a+b\otimes b\otimes b$ to $a\otimes a\otimes a+a\otimes b\otimes b$.
  2. 2.
    We need to calculate the stabilizer, which relates to the uniqueness of the corresponding rank decomposition. However, this is much more involved for tensors than for matrices. The degree of uniqueness, or identifiability, of tensor decompositions is often studied from an algebraic geometric perspective, and broader results have to assume some symmetry, small size, or low rank. See for example [Chiantini14, Blomenhofer24].
In this paper, we consider three notions of rank: canonical polyadic (CP) rank, tensor train ($T$) rank, and multilinear rank. The two obstacles are overcome by:
  1. 1.
    Restricting to a Zariski open subset where the action is transitive.
  2. 2.
    Restricting to ranks where we understand identifiability.
We thus identify three families of fixed-rank tensor manifolds as smooth homogeneous manifolds. For each of these, there is a so-called canonical Riemannian metric. We show how the geodesics in this metric can be computed efficiently.
Related work
As already mentioned, the space of matrices with fixed rank is a homogeneous manifold [Vandereycken12, Absil15, MuntheKaas15].
In low-rank approximation, the manifold perspective has been used to study several different fixed-rank tensor spaces. We mention here CP rank [Breiding18], $T$ rank [Holtz12, Uschmajew20], hierarchical Tucker rank [Uschmajew13], and multilinear rank [Koch10].
In [Uschmajew20], it is mentioned that “the set of all tensors with canonical rank bounded by $k$ is typically not closed. Moreover, while the closure of this set is an algebraic variety, its smooth part is in general not equal to the set of tensors of fixed rank $k$ and does not admit an easy explicit description.”. Our contribution is to find an “easy explicit description” for a subset of those manifolds.
We also mention that fixed-rank tensor manifolds have, to our knowledge, not previously been equipped with a Riemannian metric with known geodesics, other than in the rank 11 case [Swijsen22b, Jacobsson24].
In Riemannian optimization, line searches and gradient descents are often performed along geodesics, or along approximate geodesics, called retractions. As already mentioned, on any homogeneous manifold there is a distinguished Riemannian metric called the canonical metric. This metric has two big advantages: geodesics on a quotient can be described in terms of geodesics on its numerator; and those geodesics are always complete, meaning they can be extended to arbitrary length. Similarly, a retraction on the quotient can be induced by a retraction on the numerator. Riemannian optimization using the canonical metric has been explored, for example, on the Grassmann [Helmke07, Sato14, Boumal15, Bendokat24], Stiefel [Edelman98, Li20, Gao22], and symplectic Stiefel [Gao21, Bendokat21] manifolds.
Lie group integrators are a class of numerical integrators for ordinary differential equations (ODEs) on manifolds [Iserles00, Christiansen11, Blanes09]. They work by replacing the manifold ODE with an appropriate Lie group ODE, and identifying the underlying manifold as homogeneous informs the choice of Lie group and the design of the integrator [MuntheKaas97, MuntheKaas99a, Celledoni03, Malham08, MuntheKaas14, MuntheKaas15].
Structure of the paper
The main results are theorems˜9, 21, and 15.
Section˜2 briefly recaps the concepts from three different fields of study—homogeneous spaces, multilinear algebra, and algebraic geometry—that we use to construct our homogeneous structure. Sections˜3, 4, and 5 apply these to sets of tensors with fixed CP rank, multilinear rank, and $T$ rank respectively. The constructions are similar to each other, and the basic results are repeated without proof in sections˜4 and 5 after having been introduced in detail in section˜3.
2 Preliminaries
Let $d\geq 3$ and let $n_{1}$, …, $n_{d}$ be integers $\geq 2$. If $\mathrm{GL}(n)$ denotes the real group of $n\times n$ invertible matrices, then
$G=\mathrm{GL}(n_{1})\times\dots\times\mathrm{GL}(n_{d})$ (2.1)
has a natural action on tensor on the set of tensors $\mathbb{R}^{n_{1}}\otimes\dots\otimes\mathbb{R}^{n_{d}}$ via
$(g_{1},\dots,g_{d})\cdot(v_{1}\otimes\dots\otimes v_{d})=(g_{1}v_{1})\otimes\dots\otimes(g_{d}v_{d}).$ (2.2)
This is the Lie group action we will use to construct our manifolds.
We will write $G\curvearrowright M$ for “the action of $G$ on $M$.
2.1 Homogeneous spaces
A homogeneous space is a quotient $A/B$ of Lie groups, along with some manifold structure that is compatible with the projection $\pi\colon A\to A/B$, $a\mapsto aB$. Lee [Lee13, chapters 7 and 21] is a standard introduction to smooth homogeneous spaces and O’Neill [Oneill83, chapter 11] is a standard introduction to Riemannian homogeneous spaces.
Fix an element $T\in\mathbb{R}^{n_{1}}\otimes\dots\otimes\mathbb{R}^{n_{d}}$ and define its orbit
$G\cdot T=\set{g\cdot T\mbox{\quad s.th.\quad}g\in G}$, (2.3)
and its stabilizer (or isotropy subgroup)
$H=\set{g\in G\mbox{\quad s.th.\quad}g\cdot T=T}$. (2.4)
There is a unique smooth manifold structure on $G\cdot T=G/H$ such that, for all $f$, $f\colon G\cdot T\to\mathbb{R}$ is smooth if and only if $f\circ\pi\colon G\to\mathbb{R}$ is smooth. Then $\pi$ is called a smooth submersion.
[…]
Notably, geodesics in $G$ whose initial velocity is horizontal will keep having horizontal velocity, allowing us to define horizontal geodesics. From this also follows a one-to-one correspondence between horizontal geodesics in $G$ through $g$ and geodesics in $G\cdot T$ through $\pi(g)$ [Cheeger08, proposition 3.31]. In terms of the manifold exponential, we can write
[…]
when $X\in T_{g}G$ is horizontal. This is useful because it allows us to lift geodesics in $G\cdot T$ to geodesics in $G$.
2.2 Algebraic geometry
To derive quotients, we will need some results from algebraic geometry. Landsberg [Landsberg12] is a good reference for the algebraic perspective on tensors. Here, we just introduce a few concepts that will be useful later. An Algebraic variety is a space that is a solution set to some algebraic equation. Especially, putting an upper bound on the rank of tensors is an algebraic condition and thus yields an algebraic variety. The link between homogeneous spaces and algebraic geometry is that the orbits $G\cdot T$ are open and dense subsets of such varieties. Particularly, they are open subsets in the Zariski topology, where the closed sets are algebraic varieties.
2.3 Multilinear algebra
Tensors are high-dimensional objects, and working with them in practice often requires using some decomposition. Kolda and Bader [Kolda09] survey the CP and Tucker decompositions, and Oseledets [Oseledets11] introduces tensor trains. Graphically, tensor decompositions can be thought of as Penrose diagrams. For example, figure˜1 shows a compact SVD decomposition. The idea is that when the dimensions of the inner edges are small enough, then the decomposed tensor is cheaper to store and to operate on.
Figure 1: Penrose diagram for the compact SVD decomposition, $UDV^{\mathsf{T}}$, of an $m\times n$ rank $r$ matrix. Edges are labeled with the vector space they represent.
CP tensors
Definition 1.
The CP rank (often just rank) of a tensor $T\in\mathbb{R}^{n_{1}}\otimes\dots\otimes\mathbb{R}^{n_{d}}$ is the smallest $r$ such that $T$ is a sum of $r$ rank 11 tensors:
$T=\sum_{j=1}^{r}v_{1}^{j}\otimes\dots\otimes v_{d}^{j},$ (2.6)
where $v_{i}^{j}\in\mathbb{R}^{n_{i}}$. Such an expression is called a CP decomposition. It is illustrated in figure˜2.
Figure 2: Penrose diagram for the CP decomposition. Edges are labeled with the vector space they represent, and $V_{i}=\begin{bmatrix}v_{i}^{1}&\dots&v_{i}^{r}\end{bmatrix}$. II is the diagonal tensor $I_{i_{1}\dots i_{d}}=1$ if $i_{1}=\dots=i_{d}$ else $I_{i_{1}\dots i_{d}}=0$.
Tucker tensors
Definition 2.
The multilinear (or Tucker) rank of a tensor $T\in\mathbb{R}^{n_{1}}\otimes\dots\otimes\mathbb{R}^{n_{d}}$ is the tuple $(t_{1},\dots,t_{d})$ with
$t_{i}=\operatorname{rank}(T\colon\mathbb{R}^{n_{i}}\to\mathbb{R}^{n_{1}}\otimes\cdots\widehat{\mathbb{R}^{n_{i}}}\cdots\otimes\mathbb{R}^{n_{d}}).$ (2.7)
Proposition 3.
If $T\in\mathbb{R}^{n_{1}}\otimes\dots\otimes\mathbb{R}^{n_{d}}$ has multilinear rank $(t_{1},\dots,t_{d})$ then $T$ can be written as
$T_{k_{1}\dots k_{d}}=\sum_{\alpha_{1}=1}^{t_{1}}\cdots\sum_{\alpha_{d}=1}^{t_{d}}C_{\alpha_{1}\dots\alpha_{d}}(G_{1})_{k_{1}}{}^{\alpha_{1}}\cdots(G_{d})_{k_{d}}{}^{\alpha_{d}}$. (2.8)
Such an expression is called a Tucker decomposition. It is illustrated in figure˜3.
Proof.
By ˜2.7, we can use the matrix rank decomposition in each mode to arrive at ˜2.8. ∎
Also, conversely, a tensor admitting a Tucker decomposition with inner dimensions $t_{1}$, …, $t_{d}$ has multilinear rank at most $(t_{1},\dots,t_{d})$. If the inner dimensions and multilinear rank are equal, then the $G_{i}$ have maximal rank.
If the $G_{i}$ have maximal rank, we could without loss of generality require that they be orthogonal. Many authors do this, but the stabilizer is easier to derive if we do not.
Figure 3: Penrose diagram for the Tucker decomposition ˜2.8. Edges are labeled with the vector space they represent.
Tensor trains
Definition 4.
The $T$ rank of a tensor $T\in\mathbb{R}^{n_{1}}\otimes\dots\otimes\mathbb{R}^{n_{d}}$ is the tuple $(s_{1},\dots,s_{d-1})$ with
$s_{i}=\operatorname{rank}(T\colon\mathbb{R}^{n_{1}}\otimes\dots\otimes\mathbb{R}^{n_{i}}\to\mathbb{R}^{n_{i+1}}\otimes\dots\otimes\mathbb{R}^{n_{d}}).$ (2.9)
Proposition 5 (Oseledets [Oseledets11, Theorem 2.1]).
If $T\in\mathbb{R}^{n_{1}}\otimes\dots\otimes\mathbb{R}^{n_{d}}$ has $T$ rank $(s_{1},\dots,s_{d-1})$ then $T$ can be written as
$T_{k_{1}k_{2}\dots k_{d}}=\sum_{\alpha_{1}=1}^{s_{1}}\cdots\sum_{\alpha_{d-1}=1}^{s_{d-1}}(F_{1})_{k_{1}\alpha_{1}}(F_{2})^{\alpha_{1}}{}_{k_{2}\alpha_{2}}\cdots(F_{d})^{\alpha_{d-1}}{}_{k_{d}}$. (2.10)
Such an expression is called a $T$ decomposition. It is illustrated in figure˜4.
Figure 4: Penrose diagram for the $T$ decomposition ˜2.10. Edges are labeled with the vector space they represent.
Note also that a converse of proposition˜5 is true, that a $T$ decomposition with inner dimensions $s_{1}$, …, $s_{d-1}$ has $T$ rank at most $(s_{1},\dots,s_{d-1})$.
3 The CP manifold
Proposition 6 (Kruskal’s theorem).
Let $T\in\mathbb{R}^{n_{1}}\otimes\dots\otimes\mathbb{R}^{n_{d}}$ be CP rank $r$ with CP decomposition ˜2.6. Define $k_{i}$, for each mode $i$, so that any size $k_{i}$ subset of the factors $v_{i}^{1}$, …, $v_{i}^{r}$ is independent. If
$2r+d-1\leq\sum_{i=1}^{d}k_{i},$ (3.1)
then the decomposition is unique up to permutation of the terms and rescaling the factors.
Landsberg [Landsberg12, Theorem 12.5.3.2] formulates Kruskal’s theorem for complex vector spaces, but if a decomposition is real and complex unique then it is of course also real unique. The condition ˜3.1 is called Kruskal’s criterion. Whenever we use (Kruskal’s theorem). in the paper, we are using it via the following corollary.
Corollary 6.1.
Let $T\in\mathbb{R}^{n_{1}}\otimes\dots\otimes\mathbb{R}^{n_{d}}$ be CP rank $r$ with CP decomposition ˜2.6. Assume that, for each mode $i$, the factors $v_{i}^{1}$, …, $v_{i}^{r}$ are linearly independent. Then the decomposition is unique up to permutation of the terms and rescaling of the factors.
Note that $v_{i}^{1}$, …, $v_{i}^{r}$ are linearly independent in $\mathbb{R}^{n_{i}}$ implies $r\leq n_{i}$ for all ii.
3.1 Smooth manifold
Let $\operatorname{rank}^{-1}(r)$ denote the set of tensors with CP rank $r$. Its closure, $\overline{\operatorname{rank}^{-1}(r)}$, is an algebraic variety.
Lemma 7.
Let $r\leq n_{i}$ for all ii. Then
$G\curvearrowright\operatorname{rank}^{-1}(r)$ (3.2)
has an open dense orbit, $\Sigma_{r}$, consisting of elements with maximal multilinear rank.
Remark.
Maximal multilinear rank in $\operatorname{rank}^{-1}(r)$ is $(r,\dots,r)$.
Proof.
CP rank is preserved by the action of $G$ since
  1. 1.
    if $\sum_{j=1}^{r}v_{1}^{j}\otimes\dots\otimes v_{d}^{j}$ is a rank $r$ decomposition of $T$, then $\sum_{j=1}^{r}(g_{1}v_{1}^{j})\otimes\dots\otimes(g_{d}v_{d}^{j})$ is a rank $r$ decomposition of $g\cdot T$, so $\operatorname{rank}(T)\geq\operatorname{rank}(g\cdot T)$, and
  2. 2.
    if $\sum_{j=1}^{r}v_{1}^{j}\otimes\dots\otimes v_{d}^{j}$ is a rank $r$ decomposition of $g\cdot T$, then $\sum_{j=1}^{r}(g_{1}^{-1}v_{1}^{j})\otimes\dots\otimes(g_{d}^{-1}v_{d}^{j})$ is a rank $r$ decomposition of $T$, so $\operatorname{rank}(g\cdot T)\geq\operatorname{rank}(T)$.
Thus the action is well-defined.
Let $e_{i}^{1}$, …, $e_{i}^{n_{i}}$ be the standard basis for $\mathbb{R}^{n_{i}}$ and define
$T=\sum_{j=1}^{r}e_{1}^{j}\otimes\dots\otimes e_{d}^{j}.$ (3.3)
We now want to show that $\Sigma_{r}=G\cdot T$ is Zariski open in $\overline{\operatorname{rank}^{-1}(r)}$. It then follows that $\Sigma_{r}$ is open and dense in $\operatorname{rank}^{-1}(r)$.
Any other tensor,
$S=\sum_{j=1}^{r}f_{1}^{j}\otimes\dots\otimes f_{d}^{j},$ (3.4)
with linearly independent $f_{i}^{1}$, …, $f_{i}^{r}$ is reached by the group element $(g_{1},\dots,g_{d})$ satisfying $g_{i}e_{i}^{j}=f_{i}^{j}$. Moreover, if $f_{i}^{1}$, …, $f_{i}^{r}$ are linearly dependent for some $i$, then $S$ can’t be reached by $G$. The complement of $\Sigma_{r}$ is hence described by a Zariski closed condition, so $\Sigma_{r}$ is Zariski open.
To show that elements in $\Sigma_{r}$ have maximal multilinear rank, note that the matrix
$T_{(i)}=\sum_{j=1}^{r}(e_{1}^{j}\otimes\cdots\widehat{e_{i}^{j}}\dots\otimes e_{d}^{j})\otimes e_{i}^{j}$ (3.5)
has rank $r$ since $e_{1}^{j}\otimes\cdots\widehat{e_{i}^{j}}\cdots\otimes e_{d}^{j}$ are $r$ linearly independent vectors. Moreover, the action of $G$ preserves multilinear rank.
To show that elements outside of $\Sigma_{r}$ do not have maximal multilinear rank, note that if $f_{i}^{1}$, …, $f_{i}^{r}$ are linearly dependent for some $i$, and
$R=\sum_{j=1}^{r}f_{1}^{j}\otimes\dots\otimes f_{d}^{j},$ (3.6)
then the column space of $R_{(i)}$ is spanned by r1r-1 vectors. ∎
We now consider the subgroup of $G$ that fixes the $T$ defined in ˜$3.$ Applying (Kruskal’s theorem)., we have the following result.
Lemma 8.
The stabilizer $H$ of $G\curvearrowright\Sigma_{r}$ consists of elements of the form
$h=\begin{bmatrix}D_{1}Q&M_{1}\\ &A_{1}\end{bmatrix}\times\dots\times\begin{bmatrix}D_{d}Q&M_{d}\\ &A_{d}\end{bmatrix}$, (3.7)
where the $D_{i}$ are diagonal invertible $r\times r$ matrices such that $D_{1}\cdots D_{d}=1$, $Q$ is a permutation matrix, the $A_{i}$ are invertible $(n_{i}-r)\times(n_{i}-r)$ matrices, and the $M_{i}$ are arbitrary $r\times(n_{i}-r)$ matrices.
Combining lemmas˜7 and 8, we have the main result of this subsection.
Theorem 9.
The set of tensors with CP rank $r$ and multilinear rank $(r,\dots,r)$ is a smooth homogeneous manifold,
$\Sigma_{r}=G/H.$ (3.8)
3.2 Representatives
A point $p=gH\in\Sigma_{r}$ can be defined by specifying $g=(g_{1},\dots,g_{d})$. However, this requires specifying $n_{1}^{2}+\dots+n_{d}^{2}$ numbers, while the dimension of $\Sigma_{r}$ is only $(n_{1}+\dots+n_{d})r-(d-1)r$. To address this, we observe that a generic $g_{i}\in\mathrm{GL}(n_{i})$ can be reduced to block lower triangular form by the action of $H$. Define $h_{i}$ via
$g_{i}=\begin{bmatrix}g_{11}&g_{12}\\ g_{21}&g_{22}\end{bmatrix}=\begin{bmatrix}g_{11}&\\ g_{21}&1\end{bmatrix}\underbrace{\begin{bmatrix}1&g_{11}^{-1}g_{12}\\ &g_{22}-g_{21}g_{11}^{-1}g_{12}\end{bmatrix}}_{h_{i}^{-1}},$ (3.9)
and note that $h=(h_{1},\dots,h_{d})\in H$. Here, we see that $g_{11}$ needs to be invertible. If $g_{11}$ is not invertible, there is a permutation matrix $P$ such that $g_{i}^{\prime}=Pg_{i}$ has an invertible block $g^{\prime}_{11}$. In this way we only need to specify $g_{11}$, $g_{21}$, and $P$ for each ii.
[…]
3.3 Riemannian manifold
$G$’s algebra, $\mathfrak{g}$, consists of elements $Z=(Z_{1},\dots,Z_{d})$ where the $Z_{i}$ are arbitrary $n_{i}\times n_{i}$ matrices. We consider the Euclidean inner product on $\mathfrak{g}$, defined as
[…]
$H$’s Lie algebra, $\mathfrak{h}$, is a Lie subalgebra of $\mathfrak{g}$ and by taking the derivative of the expression ˜3.7 in lemma˜8 we see that $\mathfrak{h}$ consists of elements $Y=(Y_{1},\dots,Y_{d})$ where the $Y_{i}$ are on block form $\begin{bmatrix}Y^{i}_{11}&Y^{i}_{12}\\ &Y^{i}_{22}\end{bmatrix}$ with $Y^{i}_{11}$ diagonal such that $Y^{1}_{11}+\dots+Y^{d}_{11}=0$, and $Y^{i}_{12}$ and $Y^{i}_{22}$ arbitrary. The orthogonal complement of $\mathfrak{h}$, which we denote $\mathfrak{m}$, thus consists of elements $X=(X_{1},\dots,X_{d})$ where the $X_{i}$ are on block form $\begin{bmatrix}X^{i}_{11}&\\ X^{i}_{21}&0\end{bmatrix}$ such that $X^{i}_{11}$ has the same diagonal for every $i$ but is otherwise arbitrary, and $X^{i}_{21}$ is arbitrary.
For any $g\in G$, the coset gHgH is a submanifold of $G$. Its tangent space at $g$ is called the vertical space, and is denoted $\mathcal{V}_{g}$. The orthogonal complement to the vertical space is called horizontal space, and is denoted $\mathcal{H}_{g}$. Thus $\mathfrak{h}$ and $\mathfrak{m}$ are the vertical and horizontal spaces at 11.
[…]
Note also that the right-invariant metric on $\mathrm{GL}(n)$ is left-$\mathrm{O}(n)$-invariant. In light of our discussion in section˜3.2, this is important because permutation matrices are orthogonal and so we can work with any representative.
We now want to derive a more explicit description of the horizontal space. First, note that $\mathcal{V}_{g}=g\mathfrak{h}$, so that the vertical space consists of elements $Y=(Y_{1},\dots,Y_{d})$ where the $Y_{i}=g_{i}\begin{bmatrix}Y_{11}&Y_{12}\\ &Y_{22}\end{bmatrix}$. From now on, we suppress the $i$ superscript, but note that $Y_{11}=Y^{i}_{11}$ and the other blocks still depends on ii. Second, if we choose a block lower triangular representative as in section˜3.2, we have
[…]
$={}$ $\operatorname{tr}\left(\begin{bmatrix}X_{11}&X_{12}\\ X_{21}&X_{22}\end{bmatrix}\begin{bmatrix}g_{11}^{-1}g_{11}^{-\mathsf{T}}&-g_{11}^{-1}g_{11}^{-\mathsf{T}}g_{21}^{\mathsf{T}}\\ -g_{21}g_{11}^{-1}g_{11}^{-\mathsf{T}}&1+g_{21}g_{11}^{-1}g_{11}^{-\mathsf{T}}g_{21}^{\mathsf{T}}\end{bmatrix}\begin{bmatrix}g_{11}Y_{11}&g_{11}Y_{12}\\ g_{21}Y_{11}&g_{21}Y_{12}+Y_{22}\end{bmatrix}^{\mathsf{T}}\right)$ (3.12)
Since the $Y_{12}$ is arbitrary, $Y^{\prime}_{12}=g_{11}Y_{12}$ is also arbitrary. Similarly, $Y^{\prime}_{22}=g_{21}Y_{12}+Y_{22}$ is arbitrary. Collecting the $Y^{\prime}_{12}$ coefficients yields the condition
$-X_{11}g_{11}^{-1}g_{11}^{-\mathsf{T}}g_{21}^{\mathsf{T}}+X_{12}(1+g_{21}g_{11}^{-1}g_{11}^{-\mathsf{T}}g_{21}^{\mathsf{T}})=0.$ (3.13)
Note that $(1+g_{21}g_{11}^{-1}g_{11}^{-\mathsf{T}}g_{21}^{\mathsf{T}})$ is invertible since it is the sum of a positive definite and positive semidefinite matrix, and so it is positive definite. We can hence solve for $X_{12}$:
$X_{12}=X_{11}\Gamma_{12},$ (3.14)
where $\Gamma_{12}=g_{11}^{-1}g_{11}^{-\mathsf{T}}g_{21}^{\mathsf{T}}(1+g_{21}g_{11}^{-1}g_{11}^{-\mathsf{T}}g_{21}^{\mathsf{T}})^{-1}$. Similarly, collecting the $Y^{\prime}_{22}$ coefficients and solving for $X_{22}$ yields
$X_{22}=X_{21}\Gamma_{12}.$ (3.15)
We also note that this implies that, for each $i$,
$X_{i}={}\begin{bmatrix}X_{11}&X_{11}\Gamma_{12}\\ X_{21}&X_{21}\Gamma_{12}\end{bmatrix}={}\begin{bmatrix}X_{11}\\ X_{21}\end{bmatrix}\begin{bmatrix}1&\Gamma_{12}\end{bmatrix}$ (3.16)
is at most rank $r$.
The power series trick
Before collecting the $Y_{11}$ coefficients, we discuss an alternative expression for $\Gamma_{12}$ that will allow us to compute it more efficiently. As it is currently written, building $\Gamma_{12}$ requires inverting the $n_{i}\times n_{i}$ matrix $1+g_{21}g_{11}^{-1}g_{11}^{-\mathsf{T}}g_{21}^{\mathsf{T}}$. Note however that this inverse is an analytic function of the rank $r$ matrix $g_{21}g_{11}^{-1}g_{11}^{-\mathsf{T}}g_{21}^{\mathsf{T}}$. Its power series is
$\frac{1}{1+x}=1-x+x^{2}-x^{3}+\cdots.$ (3.17)
If $M=AB$ is a rank $r$ decomposition of an $n\times n$ matrix $M$, then
$\frac{1}{1+M}={}$ $1-AB+(AB)^{2}-(AB)^{3}+\cdots$ (3.18)
$={}$ $1+A(-1+BA-(BA)^{2}+\cdots)B$ (3.19)
$={}$ $1-A\frac{1}{1+BA}B.$ (3.20)
But BABA is an $r\times r$ matrix, allowing 1/(1+BA)-1/(1+BA) to be evaluated efficiently. This trick is also useful later for evaluating the matrix exponential. For $\Gamma_{12}$, we have the following expression,
$\Gamma_{12}={}$ $g_{11}^{-1}\left(1-\frac{\left[g_{11}^{-\mathsf{T}}g_{21}^{\mathsf{T}}g_{21}g_{11}^{-1}\right]}{1+\left[g_{11}^{-\mathsf{T}}g_{21}^{\mathsf{T}}g_{21}g_{11}^{-1}\right]}\right)g_{11}^{-\mathsf{T}}g_{21}^{\mathsf{T}}.$ (3.21)
In particular, the expression in brackets is an $r\times r$ matrix.
We are now ready to collect the $Y_{11}$ coefficients. We get the condition that the $k$th column of
$\begin{bmatrix}X_{11}\\ X_{21}\end{bmatrix}g_{11}^{-1}g_{11}^{-\mathsf{T}}-\begin{bmatrix}X_{12}\\ X_{22}\end{bmatrix}g_{21}g_{11}^{-1}g_{11}^{-\mathsf{T}}=\begin{bmatrix}X_{11}\\ X_{21}\end{bmatrix}\left(g_{11}^{-1}g_{11}^{-\mathsf{T}}-\Gamma_{12}g_{21}g_{11}^{-1}g_{11}^{-\mathsf{T}}\right)$ (3.22)
has the same dot product with the $k$th column of $\begin{bmatrix}g_{11}\\ g_{12}\end{bmatrix}$ for every ii.
This completes the description of the horizontal space $\mathcal{H}_{g}$. The most important takeaway is that horizontal vectors have a decomposition ˜3.16.
3.4 Riemannian homogeneous manifold
The construction described in section˜3.3 uses a right-invariant metric on $G$, but the resulting metric on $\Sigma_{r}$ is not necessarily right-invariant. In fact, not all smooth homogeneous manifolds allow for an invariant metric. The underlying issue is that since $h\cdot p=p$ for all $h\in H$, the inner product on $T_{p}\Sigma_{r}$ has to be invariant under $H\curvearrowright\Sigma_{r}$, but such an inner product might not exist.
The technical condition that we want to satisfy is that there exists a subspace $\mathfrak{p}\subset\mathfrak{g}$, with $\mathfrak{p}\oplus\mathfrak{h}=\mathfrak{g}$, such that $h\mathfrak{p}h^{-1}\subset\mathfrak{p}$ for all $h\in H$. $G/H$ is then called reductive and $\mathfrak{p}$ is called an invariant subspace. See O’Neill [Oneill83, Chapter 11] for more details.
Proposition 10.
$\Sigma_{r}$ is reductive if and only if $r=n_{1}=\dots=n_{d}$. Moreover, when $\Sigma_{r}$ is reductive, the $\mathfrak{m}$ that we defined in section˜3.3 is an invariant subspace.
Proposition˜10 is a higher-order version of [Vandereycken12, Proposition 3.4] and [MuntheKaas15, Proposition 5.7], which say that fixed-rank matrices are only reductive when they are square and full rank.
Proof.
Assume $r=n_{1}=\dots=n_{d}$ and let $h=D_{1}Q\times\dots\times D_{d}Q\in H$. Let $X\in\mathfrak{m}$ and consider the expression
$hXh^{-1}=(D_{1}QX_{1}Q^{-1}D_{1}^{-1},\dots,D_{d}QX_{d}Q^{-1}D_{d}^{-1}).$ (3.23)
If $\sigma$ denotes the permutation that corresponds to $Q$, then on the diagonals we have that
$(D_{i}QX_{i}Q^{-1}D_{i}^{-1})_{kk}={}$ $(D_{i})_{kk}(QX_{i}Q^{-1})_{kk}(D_{i}^{-1})_{kk}$
$={}$ $(D_{i})_{kk}(X_{i})_{\sigma(k)\sigma(k)}(D_{i}^{-1})_{kk}$
$={}$ $(X_{i})_{\sigma(k)\sigma(k)}.$ (3.24)
Hence $hXh^{-1}$ is contained in $\mathfrak{m}$, which is what we wanted to show.
Conversely, assume there exists an invariant subspace $\mathfrak{p}$ when $n_{i}>r$ for some ii. Note that elements in $\mathfrak{p}$ are determined by their projection to $\mathfrak{m}$. Concretely, elements in $\mathfrak{p}$ are determined by their $r$ first columns. So consider an element $X=(X_{1},\dots,X_{d})\in\mathfrak{p}$ with
$X_{i}=\begin{bmatrix}X_{11}&X_{12}\\ X_{21}&X_{22}\end{bmatrix}$. (3.25)
$X_{12}$ and $X_{22}$ are functions of $X_{11}$ and $X_{21}$. For our contradiction, we will show that $X_{11}$ is a multiple of the identity. Then the dimension of $\mathfrak{p}$ is too low to be complementary to $\mathfrak{h}$.
First, assume $X_{21}=0$ and let $h=\begin{bmatrix}1&M\\ &1\end{bmatrix}$ with $M$ arbitrary. Then $h^{-1}=\begin{bmatrix}1&-M\\ &1\end{bmatrix}$ and
$hX_{i}h^{-1}={}$ $\begin{bmatrix}X_{11}&-X_{11}M+X_{12}+MX_{22}\\ &X_{22}\end{bmatrix}$ (3.26)
$h^{-1}X_{i}h={}$ $\begin{bmatrix}X_{11}&X_{11}M+X_{12}-MX_{22}\\ &X_{22}\end{bmatrix}$. (3.27)
The first $r$ columns of $hX_{i}h^{-1}$ and $h^{-1}X_{i}h$ are the same, so by our previous comment they must be equal. Thus
$X_{11}M-MX_{22}=0$ (3.28)
for all $M$. Writing this as
$(X_{11}\otimes 1-1\otimes X_{22})\operatorname{vec}M=0,$ (3.29)
it is clear that the only solutions are when $X_{11}$ and $X_{22}$ are multiples of identity matrices and have the same norm.
Second, if $X_{21}\neq 0$, we can use the same argument on
$\begin{bmatrix}X_{11}&X_{12}\\ X_{21}&X_{22}\end{bmatrix}-\begin{bmatrix}&X^{\prime}_{12}\\ X_{21}&X^{\prime}_{22}\end{bmatrix}\in\mathfrak{p}$ (3.30)
to see that $X_{11}$ again is a multiple of the identity.
Whenever $r\geq 2$, the above is enough for a contradiction. However, when $r=1$, $X_{11}$ is just a $1\times 1$ matrix, and so showing that it is a multiple of the identity yields no contradiction. We need to change the argument slightly. Consider now instead the case $X_{11}=0$.
$hX_{i}h^{-1}={}$ $\begin{bmatrix}MX_{21}&*\\ X_{21}&X_{22}-X_{21}M\end{bmatrix}$. (3.31)
By the previous paragraph, we are free to add and subtract any multiple of the identity matrix and still stay in $\mathfrak{p}$. If we subtract the real number $MX_{21}$ from the diagonal, we are left with
$\begin{bmatrix}&*\\ X_{21}&X_{22}-X_{21}M-MX_{21}\cdot 1\end{bmatrix}$. (3.32)
Similarly to before, since the first column is the same as $X_{i}$, it must be equal to $X_{i}$, but for this to be true for all $M$ implies $X_{21}=0$. $X_{21}$ is ours to choose, so any restriction on it is a contradiction. ∎
3.5 Geodesics
Since the subgroup associated with $Q$ is discrete, we may for the purposes of this subsection ignore it and set $Q=1$.
Let $g_{i}\in\mathrm{GL}(n_{i})$ and $X_{i}\in T_{g_{i}}\mathrm{GL}(n_{i})$. Andruchow, Larotonda, Recht, and Varela [Andruchow14] show that the geodesics on the general linear group are [^3] Note that they use a left-invariant metric while we use a right-invariant.
$\operatorname{exp}_{g_{i}}(X_{i})=\operatorname{mexp}(X_{i}g_{i}^{-1}-(X_{i}g_{i}^{-1})^{\mathsf{T}})\operatorname{mexp}((X_{i}g_{i}^{-1})^{\mathsf{T}})g_{i}.$ (3.33)
[…]
First, recall that horizontal vectors are on the form ˜3.16. We thus have a rank $r$ decomposition
$X_{i}g_{i}^{-1}=\begin{bmatrix}X_{11}\\ X_{21}\end{bmatrix}\begin{bmatrix}g_{11}^{-1}-\Gamma_{12}g_{21}g_{11}^{-1}&\Gamma_{12}\end{bmatrix}$. (3.34)
To compute the matrix exponential of such a vector, we can use the same power series trick as before. If we name the factors in ˜3.34 $A\in\mathbb{R}^{n_{i}\times r}$ and $B\in\mathbb{R}^{r\times n_{i}}$ respectively, then
$\operatorname{mexp}(X_{i}g_{i}^{-1})={}$ $1+AB+\frac{1}{2}(AB)^{2}+\dots$
$={}$ $1+A\psi_{1}\left(BA\right)B,$ (3.35)
where $\psi_{1}(x)=1+\frac{1}{2}x+\frac{1}{3!}x^{2}+\cdots$ is the Taylor series for $(\mathrm{e}^{x}-1)/x$. The notation $\psi_{1}$ comes from the theory of exponential integrators, where the functions
$\psi_{k}(x)=\sum_{j=0}^{\infty}\frac{1}{(j+k)!}x^{j}$ (3.36)
are known as the $\psi$ functions. Importantly, the argument to $\psi_{1}$ is an $r\times r$ matrix, so it can be evaluated cheaply. We will discuss exactly how shortly.
Second, we have a rank $2r$ decomposition
$X_{i}g_{i}^{-1}-(X_{i}g_{i}^{-1})^{\mathsf{T}}=\begin{bmatrix}X_{11}&-(g_{11}^{-1}-\Gamma_{12}g_{21}g_{11}^{-1})^{\mathsf{T}}\\ X_{21}&-\Gamma_{12}^{\mathsf{T}}\end{bmatrix}\begin{bmatrix}g_{11}^{-1}-\Gamma_{12}g_{21}g_{11}^{-1}&\Gamma_{12}\\ X_{11}^{\mathsf{T}}&X_{21}^{\mathsf{T}}\end{bmatrix}$. (3.37)
So, similarly to before, denoting the factors $A^{\prime}\in\mathbb{R}^{n_{i}\times 2r}$ and $B^{\prime}\in\mathbb{R}^{2r\times n_{i}}$ respectively,
$\operatorname{mexp}(X_{i}g_{i}^{-1}-(X_{i}g_{i})^{\mathsf{T}})=1+A^{\prime}\psi_{1}(B^{\prime}A^{\prime})B^{\prime}.$ (3.38)
Here, the argument to $\psi_{1}$ is an $2r\times 2r$ matrix, so it can be evaluated cheaply.
Putting all this into ˜3.33 and doing the multiplications, we find an expression for the first $r$ columns,
$\operatorname{exp}_{g_{i}}(X_{i})\begin{bmatrix}1\\ 0\end{bmatrix}$
$=\begin{bmatrix}g_{11}\\ g_{21}\end{bmatrix}+B^{\mathsf{T}}\psi_{1}(BA)^{\mathsf{T}}A^{\mathsf{T}}\begin{bmatrix}g_{11}\\ g_{21}\end{bmatrix}+A^{\prime}\psi_{1}(B^{\prime}A^{\prime})B^{\prime}\begin{bmatrix}g_{11}\\ g_{21}\end{bmatrix}+B^{\mathsf{T}}\psi_{1}(BA)^{\mathsf{T}}A^{\mathsf{T}}A^{\prime}\psi_{1}(B^{\prime}A^{\prime})B^{\prime}\begin{bmatrix}g_{11}\\ g_{21}\end{bmatrix}$ (3.39)
Advantageously, if the multiplications are done in the right order this does not require forming any $n_{i}\times n_{i}$ matrices.
We now return to how to compute $\psi_{1}$. We will do it similarly to how the matrix exponential is usually computed, using Padé approximation and scaling and squaring. See Moler and van Loan [Moler03, methods 2 and 3] and Higham [Higham08, sections 10.3 and 10.7.4]. Let $M$ be an $r\times r$ matrix and assume first that $\norm{M}\leq 1/2$. Then $\psi_{1}(M)$ is approximated to double precision from its degree $(6,6)$ Padé approximant [Higham08, theorem 10.31]
$r_{66}(M)=\frac{1+M/26+5M^{2}/156+M^{3}/858+M^{4}/5720+M^{5}/205920+M^{6}/8648640}{1-6M/13+5M^{2}/52-5M^{3}/429+M^{4}/1144-M^{5}/25740+M^{6}/1235520}.$ (3.40)
We prove this in lemma˜24.
On the other hand, if $\norm{M}>1/2$, then we do not use the Padé approximant directly. Let $z$ be an integer such that $\norm{M}2^{-z}\leq 1/2$. Keeping in mind that these expressions are only shorthands for their Taylor series, we have
$\psi_{1}(M)={}$ $\frac{\operatorname{mexp}(M)-1}{M}$
$={}$ $\frac{\operatorname{mexp}(2^{-z}M)^{2^{z}}-1}{M}$
$={}$ $\frac{\operatorname{mexp}(2^{-z}M)-1}{M}(\operatorname{mexp}(2^{-z}M)+1)\cdots(\operatorname{mexp}(2^{-1}M)+1)$
$={}$ $2^{-z}\psi_{1}(2^{-z}M)(\operatorname{mexp}(2^{-z}M)+1)\cdots(\operatorname{mexp}(2^{-1}M)+1).$ (3.41)
In the second to last step, we used $x^{2}-1=(x-1)(x+1)$ recursively. This scaling and squaring step is similar to the algorithm proposed by Hochbruck, Lubich, and Selhofer [Hochbruck98].
Counting the operations
[…]
Proposition 11.
Given a tangent vector $X\in T_{p}\Sigma_{r}$, $\operatorname{exp}_{p}(X)$ can be estimated [^4] Meaning that everything is computed exactly except for $\psi_{1}$, which is Padé approximated to within double precision. using
[…]
where [^5] This formula is valid for any norm, as long as it is the same as in appendix A. $z_{i}=\lceil\operatorname{log}_{2}\norm{X_{i}g_{i}^{-1}}\rceil+2$.
See appendix˜B for the proof.
This can be viewed as a tensorial version of Vandereycken’s et al. [Vandereycken12] corollary 4.2. We also mention that the effects of rounding errors in the Padé approximant and squaring step are not completely understood [Moler03], and so we do not do a full error analysis.
4 The Tucker manifold
In proposition˜3, we used the matrix rank decomposition recursively, and since that decomposition is unique up to a change of basis in the inner vector space, we immedeately arrive at the following.
Proposition 12.
The Tucker decomposition is unique up to a change of basis in $\mathbb{R}^{t_{1}}$, …, $\mathbb{R}^{t_{d}}$. More precisely,
$\sum_{\alpha_{1}=1}^{t_{1}}\cdots\sum_{\alpha_{d}=1}^{t_{d}}C^{\prime}_{\alpha_{1}\dots\alpha_{d}}(G^{\prime}_{1})_{k_{1}}{}^{\alpha_{1}}\cdots(G^{\prime}_{d})_{k_{d}}{}^{\alpha_{d}}=\sum_{\alpha_{1}=1}^{t_{1}}\cdots\sum_{\alpha_{d}=1}^{t_{d}}C_{\alpha_{1}\dots\alpha_{d}}(G_{1})_{k_{1}}{}^{\alpha_{1}}\cdots(G_{d})_{k_{d}}{}^{\alpha_{d}}$ (4.1)
iff there are matrices $U_{1}\in\mathrm{GL}(t_{1})$, …, $U_{d}\in\mathrm{GL}(t_{d})$ such that
$(G^{\prime}_{i})_{k_{i}}{}^{\alpha_{i}}$={} $\sum_{\beta=1}^{t_{i}}(G_{i})_{k_{i}}{}^{\beta}(U_{i})_{\beta}{}^{\alpha_{i}},\quad i=1,\dots,d$ (4.2)
$(C^{\prime})_{\alpha_{1}\dots\alpha_{d}}={}$ $\sum_{\beta_{1}=1}^{t_{1}}\cdots\sum_{\beta_{d}=1}^{t_{d}}C_{\beta_{1}\dots\beta_{d}}(U_{1}^{-1})^{\beta_{1}}{}_{\alpha_{1}}\cdots(U_{d}^{-1})^{\beta_{d}}{}_{\alpha_{d}}$. (4.3)
Many of the arguments in this section are the same as in section˜3, so we do not repeat them here but just refer back.
4.1 Smooth manifold
Let $\operatorname{mrank}^{-1}(t_{1},\dots,t_{d})$ denote the set of tensors with multilinear rank $(t_{1},\dots,t_{d})$. Like $\operatorname{ttrank}^{-1}(s_{1},\dots,s_{d-1})$ and $\operatorname{rank}^{-1}(r)$, its closure is an algebraic variety.
Lemma 13.
Let $t_{1}=t_{2}\cdots t_{d}$. Then
$G\curvearrowright\Lambda_{t_{1}\dots t_{d}}:=\operatorname{mrank}^{-1}(t_{1},\dots,t_{d})$ (4.4)
is a transitive action.
Proof.
Let $E_{i}$ be the identity matrix $I_{n_{i}\times t_{i}}$ seen as an element of $\mathbb{R}^{n_{i}}\otimes\mathbb{R}^{t_{i}}$ and let II be the identity matrix $I_{t_{1}\times t_{1}}$ seen as an element of $\mathbb{R}^{t_{1}}\otimes\dots\otimes\mathbb{R}^{t_{d}}$. Then define
$T_{k_{1}\dots k_{d}}=\sum_{\alpha_{1},\dots,\alpha_{d}}I_{\alpha_{1}\dots\alpha_{d}}(E_{1})_{k_{1}}{}^{\alpha_{1}}\cdots(E_{d})_{k_{d}}{}^{\alpha_{d}}$. (4.5)
Now, fix $C$ and $G_{1}$, …, $G_{i}$. By a previous comment, a Tucker decomposition with inner dimensions $t_{1}$, …, $t_{d}$ has $G_{i}$ with linearly independent columns. So let $\overline{G}_{i}\colon\mathbb{R}^{n_{i}\times(n_{i}-t_{i})}$ be a basis completion to $G_{i}$. Moreover, the unfolding $C\colon\mathbb{R}^{t_{1}}\to\mathbb{R}^{t_{2}}\otimes\dots\otimes\mathbb{R}^{t_{d}}$ is an invertible $t_{1}\times t_{1}$ matrix. The tensor
$S_{k_{1}\dots k_{d}}=\sum_{\alpha_{1},\dots,\alpha_{d}}C_{\alpha_{1}\dots\alpha_{d}}(G_{1})_{k_{1}}{}^{\alpha_{1}}\cdots(G_{d})_{k_{d}}{}^{\alpha_{d}}$. (4.6)
[…]
There are other cases than $t_{1}=t_{2}\cdots t_{d}$ where $\operatorname{mrank}^{-1}(t_{1},\dots,t_{d})$ has an open and dense orbit. However, describing those orbits is more work and involves the so-called castling transform of the factors, which generalizes the statement that $\mathbb{R}^{p}\otimes\mathbb{R}^{q}\otimes\mathbb{R}^{r}$ has an open orbit iff $\mathbb{R}^{p}\otimes\mathbb{R}^{q}\otimes\mathbb{R}^{pq-r}$ has an open orbit. See Venturelli [Venturelli18] or Landsberg [Landsberg12, Section 10.2.2] for details. We also repeat the observation from the introduction: that there is typically more degrees of freedom in $\mathbb{R}^{t_{1}}\otimes\dots\otimes\mathbb{R}^{t_{d}}$, namely $t_{1}\cdots t_{d}$ degrees, than in $\mathrm{GL}(t_{1})\times\dots\times\mathrm{GL}(t_{d})$, namely $t_{1}^{2}+\dots+t_{d}^{2}$ degrees. This imposes the restriction
$t_{1}\cdots t_{d}\leq t_{1}^{2}+\dots+t_{d}^{2}.$ (4.7)
Furthermore, since $t_{1}>t_{2}\cdots t_{d}$ is not a possible multilinear rank, we also have the restriction
$t_{1}\leq t_{2}\cdots t_{d}.$ (4.8)
Solving for $t_{1}$, we find that it must satisfy
$\frac{t_{2}\cdots t_{d}+\sqrt{(t_{2}\cdots t_{d})^{2}-4(t_{2}^{2}+\dots+t_{d}^{2})}}{2}\leq{}$ $t_{1}\leq t_{2}\cdots t_{d}.$ (4.9)
This domain is typically very small. For $d=3$, $t_{2}=t_{3}=10$ for example, we have $98\leq t_{1}\leq 100$. We therefore settle for describing the case $t_{1}=t_{2}\dots t_{d}$.
From proposition˜12 we can directly compute the stabilizer.
Lemma 14.
The stabilizer $H$ of $G\curvearrowright\Lambda_{t_{1}\dots t_{d}}$ consists of elements of the form
$\begin{bmatrix}A_{2}^{-\mathsf{T}}\otimes\dots\otimes A_{d}^{-\mathsf{T}}&M_{1}\\ &B_{1}\end{bmatrix}\times\begin{bmatrix}A_{2}&M_{2}\\ &B_{2}\end{bmatrix}\times\dots\times\begin{bmatrix}A_{d}&M_{d}\\ &B_{d}\end{bmatrix}$ (4.10)
where the $A_{i}$ are invertible $t_{i}\times t_{i}$ matrices, the $B_{i}$ are invertible $(n_{i}-t_{i})\times(n_{i}-t_{i})$ matrices, and $M_{i}$ are $t_{i}\times(n_{i}-t_{i})$ matrices.
Theorem 15.
The set of tensors with multilinear rank $(t_{1},\dots,t_{d})$, where $t_{1}=t_{2}\cdots t_{d}$, is a smooth homogeneous manifold,
$\Lambda_{t_{1}\dots t_{d}}=G/H.$ (4.11)
4.2 Representatives
Similarly to section˜3.2, if $g=(g_{1},\dots,g_{d})\in G$, then we only need to store the first $t_{i}$ columns of $g_{i}$.
4.3 Riemannian manifold
Similarly to section˜3.3, we can take the derivative of the expression ˜4.10 in lemma˜14 to get $H$’s Lie algebra, $\mathfrak{h}$. It consists of elements $Y=(Y_{1},\dots,Y_{d})$ where the $Y_{i}$ are on block form
$Y_{1}={}$ $\begin{bmatrix}-K_{2}^{\mathsf{T}}\otimes 1\otimes\dots\otimes 1-\dots-1\otimes\dots\otimes 1\otimes K_{d}^{\mathsf{T}}&*\\ &*\end{bmatrix}$,
$Y_{i}={}$ $\begin{bmatrix}K_{i}&*\\ &*\end{bmatrix},\quad 2\leq i\leq d,$ (4.12)
for arbitrary $t_{i}\times t_{i}$ matrices $K_{i}$.
$\mathfrak{h}$’s orthogonal complement, $\mathfrak{m}$, consists of elements $X=(X_{1},\dots,X_{d})$ where the $X_{i}$ are on block form
$X_{1}={}$ $\begin{bmatrix}L_{1}&\\ *&0\end{bmatrix}$,
$X_{i}={}$ $\begin{bmatrix}L_{i}&\\ *&0\end{bmatrix},\quad 2\leq i\leq d,$ (4.13)
such that
$L_{i}={}$ $\operatorname{tr}_{i}L_{1},\quad 2\leq i\leq d,$ (4.14)
where $\operatorname{tr}_{i}(A_{1}\otimes\dots\otimes A_{d})=(\operatorname{tr}A_{1})\cdots\widehat{\operatorname{tr}A_{i}}\cdots(\operatorname{tr}A_{d})A_{i}^{\mathsf{T}}$. The easiest way to see this restriction on $L_{i}$ is to write the inner product as
[…]
and note that this should hold for all $K_{i}$ separately. Note that $L_{2}$, …, $L_{d}$ are completely determined by $L_{1}$.
Like in section˜3.3, we consider the right-invariant metric on $G$ induced by the Euclidean inner product on $\mathfrak{g}$, and define the metric on $\Lambda_{t_{1}\dots t_{d}}$ by demanding that $\pi\colon G\to\Lambda_{t_{1}\dots t_{d}}$ is a Riemannian submersion. We do not give a full description of the horizontal space at a general point, but just note that, for each $i$, the same argument that was used in section˜3.3 can be used to derive a rank $t_{i}$ decomposition similar to ˜3.16.
4.4 Riemannian homogeneous manifold
Proposition 16.
$\Lambda_{t_{1}\dots t_{d}}$ is reductive if and only if $n_{1}=t_{1}$, …, $n_{d}=t_{d}$. Moreover, when $\Lambda_{t_{1}\dots t_{d}}$ is reductive, $\mathfrak{m}$ is an invariant subspace.
Proof.
Assume $n_{1}=t_{1}$, …, $n_{d}=t_{d}$ and let $h=((A_{2}\otimes\dots\otimes A_{d})^{\mathsf{T}},A_{2},\dots,A_{d})\in H$. We have to show that ˜4.14 is preserved by $L_{i}\mapsto L_{i}^{\prime}=A_{i}L_{i}A_{i}^{-1}$. This can be seen from
$L_{i}^{\prime}={}$ $A_{i}L_{i}A_{i}^{-1}$
$={}$ $A_{i}(\operatorname{tr}_{i}L_{1})A_{i}^{-1}$
$={}$ $\operatorname{tr}_{i}[(A_{1}\otimes\dots\otimes A_{d})^{-\mathsf{T}}L_{1}(A_{1}\otimes\dots\otimes A_{d})^{\mathsf{T}}]$
$={}$ $\operatorname{tr}_{i}L_{1}^{\prime}.$ (4.16)
If $n_{i}\neq t_{i}$ for some $i$, then it is possible to show that $\Lambda_{t_{1}\dots t_{d}}$ is not reductive with the same argument as in proposition˜10. ∎
4.5 Geodesics
By an argument completely analogous to proposition˜11, we have the following result.
Proposition 17.
Given a tangent vector $X\in T_{p}\Lambda_{t_{1}\dots t_{d}}$, $\operatorname{exp}_{p}(X)$ can be estimated using
[…]
where $z_{i}=\lceil\operatorname{log}_{2}\norm{X_{i}g_{i}^{-1}}\rceil+2$.
5 The tensor train manifold
Proposition 18.
The $T$ decomposition is unique up to a change of basis in $\mathbb{R}^{s_{1}}$, …, $\mathbb{R}^{s_{d-1}}$. More precisely,
$\sum_{\alpha_{1}}^{s_{1}}\cdots\sum_{\alpha_{d-1}}^{s_{d-1}}(F_{1}^{\prime})_{k_{1}\alpha_{1}}(F_{2}^{\prime})^{\alpha_{1}}{}_{k_{2}\alpha_{2}}\cdots(F_{d}^{\prime})^{\alpha_{d-1}}{}_{k_{d}}=\sum_{\alpha_{1}}^{s_{1}}\cdots\sum_{\alpha_{d-1}}^{s_{d-1}}(F_{1})_{k_{1}\alpha_{1}}(F_{2})^{\alpha_{1}}{}_{k_{2}\alpha_{2}}\cdots(F_{d})^{\alpha_{d-1}}{}_{k_{d}}$ (5.1)
iff there are matrices $U_{1}\in\mathrm{GL}(s_{1})$, …, $U_{d-1}\in\mathrm{GL}(s_{d-1})$ such that
$(F_{1}^{\prime})_{k_{i}\alpha_{i}}={}$ $\sum_{\beta=1}^{s_{1}}(F_{1})_{k_{1}\beta}(U_{2}^{-1})^{\beta}{}_{\alpha_{1}}$, (5.2)
$(F_{i}^{\prime})^{\alpha_{i-1}}{}_{k_{i}\alpha_{i}}$={} $\sum_{\beta=1}^{s_{i-1}}\sum_{\gamma=1}^{s_{i}}(U_{i-1})_{\beta}{}^{\alpha_{i-1}}(F_{i})^{\beta}{}_{k_{i}\gamma}(U_{i}^{-1})^{\gamma}{}_{\alpha_{i}},\quad 2\leq i\leq d-1,$ (5.3)
$(F_{d}^{\prime})^{\alpha_{d-1}}{}_{k_{d}}$={} $\sum_{\beta=1}^{s_{d-1}}(U_{d-1})_{\beta}{}^{\alpha_{d-1}}(F_{d})^{\beta}{}_{k_{d}}$. (5.4)
Uschmajew and Vandereycken [Uschmajew13, proposition 3] shows that this holds for Hierarchical Tucker decompositions, of which the $T$ decomposition is a special case.
5.1 Smooth manifold
Let $\operatorname{ttrank}^{-1}(s_{1},\dots,s_{d-1})$ denote the set of tensors with $T$ rank $(s_{1},\dots,s_{d-1})$. Like $\operatorname{rank}^{-1}(r)$, its closure is an algebraic variety.
Lemma 19.
Let $s_{1}\leq n_{1}$, $s_{i-1}s_{i}\leq n_{i}$ for all $2\leq i\leq d-1$, and $s_{d-1}\leq n_{d}$. Then
$G\curvearrowright\operatorname{ttrank}^{-1}(s_{1},\dots,s_{d-1})$ (5.5)
has an open dense orbit, $\Pi_{s_{1}\dots s_{d-1}}$, consisting of elements with maximal multilinear rank.
Remark.
Maximal multilinear rank in $\operatorname{ttrank}^{-1}(s_{1},\dots,s_{d-1})$ is $(s_{1},s_{1}s_{2},\dots,s_{d-1})$.
Proof.
From ˜2.9, it is clear that $T$ rank is preserved under $G$. The action is thus well-defined.
Let $E_{i}\colon\mathbb{R}^{s_{i-1}}\otimes\mathbb{R}^{n_{i}}\otimes\mathbb{R}^{s_{i}}$ be the tensor whose unfolding is the identity matrix, $(E_{i}\colon\mathbb{R}^{n_{i}}\to\mathbb{R}^{s_{i-1}}\otimes\mathbb{R}^{s_{i}})=I_{n_{i}\times s_{i-1}s_{i}}$, and define
$T_{k_{1}k_{2}\dots k_{d}}=\sum_{\alpha_{1},\dots,\alpha_{d-1}}(E_{1})_{k_{1}\alpha_{1}}(E_{2})^{\alpha_{1}}{}_{k_{2}\alpha_{2}}\cdots(E_{d})^{\alpha_{d-1}}{}_{k_{d}}$. (5.6)
Similarly to the proof of lemma˜7, we now want to show that $\Pi_{s_{1}\dots s_{d-1}}=G\cdot T$ is Zariski open in $\overline{\operatorname{ttrank}^{-1}(s_{1},\dots,s_{d-1})}$.
Any other tensor
$S_{k_{1}\dots k_{d}}=\sum_{\alpha_{1},\dots,\alpha_{d-1}}(F_{1})_{k_{1}\alpha_{1}}(F_{2})^{\alpha_{1}}{}_{k_{2}\alpha_{2}}\cdots(F_{d})^{\alpha_{d-1}}{}_{k_{d}}$ (5.7)
where the matrices $F_{i}\colon\mathbb{R}^{n_{i}\times s_{i-1}s_{i}}$ have full rank is reached by the group element $(g_{1},\dots,g_{d})$ where $g_{i}=\begin{bmatrix}F_{i}&\overline{F}_{i}\end{bmatrix}$ such that $\overline{F}_{i}\colon\mathbb{R}^{n_{i}\times(n_{i}-s_{i-1}s_{i})}$ is a basis completion. Moreover, the action of $G$ preserves the rank of $F_{i}$. The condition that the $F_{i}$ have full rank is a Zariski open condition.
To show that elements outside of $\Pi_{s_{1}\dots s_{d-1}}$ don not have maximal multilinear rank, note that the multilinear rank of $S$ is just $(\operatorname{rank}F_{1},\dots,\operatorname{rank}F_{d})$. ∎
From Proposition˜18 we can directly compute the stabilizer.
Lemma 20.
The stabilizer $H$ of $G\curvearrowright\Pi_{s_{1}\dots s_{d-1}}$ consists of elements of the form
$\begin{bmatrix}A_{1}&M_{1}\\ &B_{1}\end{bmatrix}\times\begin{bmatrix}A_{1}^{-\mathsf{T}}\otimes A_{2}&M_{2}\\ &B_{2}\end{bmatrix}\times\dots\times\begin{bmatrix}A_{d-1}^{-\mathsf{T}}&M_{d}\\ &B_{d}\end{bmatrix}$ (5.8)
where the $A_{i}$ are invertible $s_{i}\times s_{i}$ matrices, the $B_{i}$ are invertible $(n_{i}-s_{i-1}s_{i})\times(n_{i}-s_{i-1}s_{i})$ matrices, and the $M_{i}$ are $s_{i-1}s_{i}\times(n_{i}-s_{i-1}s_{i})$ matrices.
Combining lemmas˜19 and 20, we have the main result of this subsection.
Theorem 21.
The set of tensors with $T$ rank $(s_{1},\dots,s_{d-1})$ and multilinear rank $(s_{1},$ $s_{1}s_{2},\dots,s_{d-1})$ is a smooth homogeneous manifold,
$\Pi_{s_{1}\dots s_{d-1}}=G/H.$ (5.9)
5.2 Representatives
Similarly to section˜3.2, if $g=(g_{1},\dots,g_{d})\in G$, then we only need to store the first $s_{i-1}s_{i}$ columns of $g_{i}$.
5.3 Riemannian manifold
Similarly to section˜3.3, we can take the derivative of the expression ˜5.8 in lemma˜20 to get $H$’s Lie algebra, $\mathfrak{h}$. It consists of elements $Y=(Y_{1},\dots,Y_{d})$ where the $Y_{i}$ are on block form
$Y_{1}={}$ $\begin{bmatrix}K_{1}&*\\ &*\end{bmatrix}$,
$Y_{i}={}$ $\begin{bmatrix}-K_{i-1}^{\mathsf{T}}\otimes 1+1\otimes K_{i}&*\\ &*\end{bmatrix},\quad 2\leq i\leq d-1,$
$Y_{d}={}$ $\begin{bmatrix}-K_{d-1}^{\mathsf{T}}&*\\ &*\end{bmatrix}$, (5.10)
for arbitrary $s_{i}\times s_{i}$ matrices $K_{i}$. The orthogonal complement, $\mathfrak{m}$, consists of elements $X=(X_{1},\dots,X_{d})$ where the $X_{i}$ are on block form
$X_{1}={}$ $\begin{bmatrix}L_{1}&\\ *&0\end{bmatrix}$,
$X_{i}={}$ $\begin{bmatrix}L_{i}&\\ *&0\end{bmatrix},\quad 2\leq i\leq d-1,$
$X_{d}={}$ $\begin{bmatrix}L_{d}&\\ *&0\end{bmatrix}$, (5.11)
such that
$L_{1}={}$ $\operatorname{tr}_{2}L_{2}$
$\operatorname{tr}_{1}L_{i-1}={}$ $\operatorname{tr}_{2}L_{i},\quad 2\leq i\leq d-1$
$\operatorname{tr}_{1}L_{d-1}={}$ $L_{d},$ (5.12)
where $\operatorname{tr}_{1}(A\otimes B)=(\operatorname{tr}A)B$ and $\operatorname{tr}_{2}(A\otimes B)=(\operatorname{tr}B)A^{\mathsf{T}}$. This can be shown the in the same way as ˜4.14.
Like in section˜3.3, we consider the right-invariant metric on $G$ induced by the Euclidean inner product on $\mathfrak{g}$, and define the metric on $\Pi_{s_{1}\dots s_{d-1}}$ by demanding that $\pi\colon G\to\Pi_{s_{1}\dots s_{d-1}}$ is a Riemannian submersion. We do not give a full description of the horizontal space at a general point $g$, but just note that the same argument that was used in section˜3.3 can be used to derive a rank $s_{i-1}s_{i}$ decomposition similar to ˜3.16,
5.4 Riemannian homogeneous manifold
Proposition 22.
$\Pi_{s_{1}\dots s_{d-1}}$ is reductive if and only if $n_{1}=s_{1}$, $n_{2}=s_{1}s_{2}$, …, $n_{d}=s_{d-1}$. Moreover, when $\Pi_{s_{1}\dots s_{d-1}}$ is reductive, $\mathfrak{m}$ is an invariant subspace.
Proof.
Assume $n_{1}=s_{1}$, $n_{2}=s_{1}s_{2}$, …, $n_{d}=s_{d-1}$ and let $h=(A_{1},A_{1}^{-\mathsf{T}}\otimes A_{2},\dots,A_{d-1}^{-\mathsf{T}})\in H$. We have that $X\mapsto hXh^{-1}$ maps $L_{i}\mapsto L_{i}^{\prime}=(A_{i-1}^{-\mathsf{T}}\otimes A_{i})L_{i}(A_{i-1}^{\mathsf{T}}\otimes A_{i}^{-1})$. To show that $\mathfrak{m}$ is invariant, we need to argue that the relation ˜5.12 is preserved. This can be seen from
$\operatorname{tr}_{1}L_{i-1}^{\prime}={}$ $\operatorname{tr}_{1}[(A_{i-2}^{-\mathsf{T}}\otimes A_{i-1})L_{i-1}(A_{i-2}^{\mathsf{T}}\otimes A_{i-1}^{-1})]$
$={}$ $A_{i-1}(\operatorname{tr}_{1}L_{i-1})A_{i-1}^{-1}$
$={}$ $A_{i-1}(\operatorname{tr}_{2}L_{i})A_{i-1}^{-1}$
$={}$ $\operatorname{tr}_{2}[(A_{i-1}^{-\mathsf{T}}\otimes A_{i})L_{i}(A_{i-1}^{\mathsf{T}}\otimes A_{i}^{-1})]$
$={}$ $\operatorname{tr}_{2}L_{i}^{\prime}.$ (5.13)
On the other hand, if $n_{i}\neq s_{i-1}s_{i}$ for some $i$, then it is possible to show that $\Pi_{s_{1}\dots s_{d-1}}$ is not reductive with the same argument as in proposition˜10. ∎
5.5 Geodesics
By an argument completely analogous to proposition˜11, we have the following result.
Proposition 23.
Given a tangent vector $X\in T_{p}\Pi_{s_{1}\dots s_{d-1}}$, $\operatorname{exp}_{p}(X)$ can be estimated using
[…]
where $z_{i}=\lceil\operatorname{log}_{2}\norm{X_{i}g_{i}^{-1}}\rceil+2$.
Appendix A Error bound for Padé approximant
[…]
Lemma 24.
If $\norm{M}\leq 1/2$, then $r_{66}(M)$ approximates $\psi_{1}(M)$ to within double precision $2^{-53}\approx${10}^{-16}$\$.
Proof.
$r_{66}(M)$ has a Taylor expansion $\sum_{i=1}^{\infty}\frac{r_{66}^{(n)}(0)}{n!}M^{n}$. Since $r_{66}$ is the Padé approximant to $\psi_{1}(M)=\sum_{i=1}^{\infty}\frac{1}{(n+1)!}M^{n}$, they agree up to the first $12$ terms. We thus want to bound the series
$\psi_{1}(M)-r_{66}(M)=\sum_{n=13}^{\infty}\frac{1-(n+1)r_{66}^{(n)}(0)}{(n+1)!}M^{n}.$ (A.1)
First, we compute term $13$ and $14$ manually. They are $M^{13}/149597947699200$ and $181M^{14}/29171599801344000$ respectively.
Second, we note that the $n$th derivative of $r_{66}$ is a rational function $p_{n}(M)/q_{n}(M)$. The quotient rule gives the recurrence relation $p_{n+1}=p_{n}^{\prime}q_{n}-p_{n}q_{n}^{\prime}$ and $q_{n+1}=q_{n}^{2}$. We have that $q_{n}(0)=1$ for all $n$. Moreover, $\norm{p_{0}^{(k)}(0)}$, $\norm{q_{0}^{(k)}(0)}\leq 1/2^{k}$ for all $k$. Using the triangle inequality in the recurrence relation, this implies that $\norm{r_{66}^{(n)}(0)}=\norm{p_{n}(0)}\leq 1$. Thus
$\norm{\psi_{1}(M)-r_{66}(M)}={}$ $\norm{\sum_{n=13}^{\infty}\frac{1-(n+1)r_{66}^{(n)}(0)}{(n+1)!}M^{n}}$
$\leq{}$ $\frac{\norm{M}^{13}}{149597947699200}+\frac{181\norm{M}^{14}}{29171599801344000}+\sum_{n=15}^{\infty}\frac{1+(n+1)\norm{r_{66}^{(n)}(0)}}{(n+1)!}\norm{M}^{n}$
$\leq{}$ $\frac{(1/2)^{13}}{149597947699200}+\frac{181(1/2)^{14}}{29171599801344000}+\sum_{n=15}^{\infty}\frac{1+(n+1)}{(n+1)!}(1/2)^{n}$
$={}$ $3\sqrt{e}-\frac{2364006584786656554317}{477947491145220096000}$
$\leq{}$ $$2.7\text{\times}{10}^{-17}$\$. (A.2)
Lemma 25.
If $\norm{M}\leq 1/2$, then $\operatorname{mexp}(M)$ is approximated by its Padé approximant to within double precision.
We do not prove lemma˜25 here, but instead refer to Higham [Higham08, Section 10.3].
Appendix B Appendix: Counting the operations
We now prove proposition˜11.
Proof.
  • $20n_{i}r^{2}/3+22r^{3}/3$ operations to build $\Gamma_{12}$ using ˜3.21:
    • $8n_{i}r^{2}/3$ operations to divide $g_{21}$ by $g_{11}$,
    • $2n_{i}r^{2}$ operations to multiply $g_{11}^{-\mathsf{T}}g_{21}^{\mathsf{T}}$ with $g_{21}g_{11}^{-1}$,
    • $2r^{3}$ operations to square $g_{11}^{-\mathsf{T}}g_{21}^{\mathsf{T}}g_{21}g_{11}^{-1}$,
    • $8r^{3}/3$ operations to divide by $1+g_{11}^{-\mathsf{T}}g_{21}^{\mathsf{T}}g_{21}g_{11}^{-1}$,
    • $8r^{3}/3$ operations to divide outside the parenthesis from the left by $g_{11}$,
    • $2n_{i}r^{2}$ to operations to multiply from the right by $g_{11}^{-\mathsf{T}}g_{21}^{\mathsf{T}}$,
  • no extra operations to build AA,
  • $2n_{i}r^{2}+8r^{3}/3$ operations to build $B$:
    • $2n_{i}r^{2}$ operations to multiply $\Gamma_{12}$ with $g_{21}$,
    • $8r^{3}/3$ operations to divide $1-\Gamma_{12}g_{21}$ by $g_{11}$ from the right,
  • $2n_{i}r^{2}$ operations to multiply $B$ and AA,
  • $(52/3+4(z_{i}-1))r^{3}$ operations to evaluate $\psi_{1}$ on BABA using ˜3.40 and 3.41,
    • $12r^{3}$ operations to form the powers $2^{-z}M$, …, $(2^{-z}M)^{6}$,
    • $8r^{3}/3$ operations to evaluate the quotient,
    • $8r^{3}/3$ operations to evaluate $\operatorname{mexp}(2^{-z}M)$,
    • $2(z-1)r^{3}$ operations to form $\operatorname{mexp}(2^{-z+1}M)$, …, $\operatorname{mexp}(2^{-1}M)$ by recursively using $\operatorname{mexp}(2M)=\operatorname{mexp}(M)\operatorname{mexp}(M)$,
    • $2(z-1)r^{3}$ operations to multiply the factors in ˜3.41,
  • no extra operations to build $A^{\prime}$ or $B^{\prime}$,
  • $2n_{i}(2r)^{2}$ operations to multiply $B^{\prime}$ and $A^{\prime}$,
  • $(52/3+4(z_{i}-1))(2r)^{3}$ operations to evaluate $\psi_{1}$ on $B^{\prime}A^{\prime}$,
  • no extra operations to build the first term in ˜3.39,
  • $4n_{i}r^{2}+2r^{3}$ operations to build the second term in ˜3.39:
    • $2n_{i}r^{2}$ operations to multiply $B$ with $\begin{bmatrix}g_{11}\\ g_{21}\end{bmatrix}$,
    • $2r^{3}$ operations to multiply $\psi_{1}(BA)$ with $B\begin{bmatrix}g_{11}\\ g_{21}\end{bmatrix}$,
    • $2n_{i}r^{2}$ operations to multiply with AA from the left,
  • $8n_{i}r^{2}+8r^{3}$ operations to build the third term in ˜3.39:
    • $2n_{i}(2r)r$ operations to multiply $B^{\prime}$ with $\begin{bmatrix}g_{11}\\ g_{21}\end{bmatrix}$,
    • $2(2r)^{2}r$ operations to multiply $\psi_{1}(B^{\prime}A^{\prime})$ with $B^{\prime}\begin{bmatrix}g_{11}\\ g_{21}\end{bmatrix}$,
    • $2n_{i}(2r)r$ operations to multiply with $A^{\prime}$ from the left,
  • $6n_{i}r^{2}+6r^{3}$ operations to build the last term in ˜3.39:
    • $2n_{i}r(2r)$ operations to multiply $B$ with $A^{\prime}$,
    • $2r(2r)r$ operations to multiply $BA^{\prime}$ with $\psi_{1}(B^{\prime}A^{\prime})B^{\prime}\begin{bmatrix}g_{11}\\ g_{21}\end{bmatrix}$,
    • $2r^{3}$ operations to multiply with $\psi_{1}(BA)$ from the left,
    • $2n_{i}r^{2}$ operations to multiply with AA from the left.